# Evaluating Foundation Models’ 3D Understanding Through Multi-View Correspondence Analysis

### v1

@@ Abstract
Benchmarking 3D spatial understanding of foundation models is essential for real-world applications such as robotics and autonomous driving. Existing evaluations often rely on downstream finetuning with linear heads or task-specific decoders, making it difficult to isolate the intrinsic 3D reasoning ability of pretrained encoders. In this work, we introduce a novel benchmark for in-context 3D scene understanding that requires no finetuning and directly probes the quality of dense visual features. Building on the Hummingbird framework, which evaluates in-context 2D scene understanding, we extend the setup to the 3D Multi-View ImageNet (MVImgNet) dataset. Given a set of images from objects in specific angles (keys), we benchmark the performance of segmenting novel views (queries) and report the scores in 4 categories of easy, medium, hard, and extreme based on the key-query view contrast. We benchmark 8 state-of-the-art foundation models and show DINO-based encoders remain competitive across large viewpoint shifts, while 3D-aware models like VGGT require dedicated multi-view adjustments. Our code is publicly available at https://github.com/ToyeshC/open-hummingbird-3d-eval.
@@ 1 Introduction
In recent years, we have seen the rise of foundation models [opportunities_risks_FoMo, llama, gpt4, qwen3]. These large-scale pre-trained models support many tasks, including visual understanding, which enables direct comparison of their encoders. Despite their impressive capabilities, state-of-the-art Vision Transformers (ViTs) can suffer catastrophic failures when objects are viewed from unusual angles, as demonstrated by recent angle-sensitivity studies [ruan2023viewpointinvariantvisualrecognitionadversarial]. However, most vision benchmarks emphasize single-view tasks or dense synthesis [mildenhall2020nerfrepresentingscenesneural], leaving segmentation robustness under camera rotations comparatively underexplored.
To address viewpoint variability, recent work has explored in-context learning (ICL) as a means to adapt models to unseen views without retraining. ICL is the ability of a model to perform new tasks by conditioning on a prompt [explanation_in-context_learning]. An example is Hummingbird [hummingbird], a memory-augmented ViT for ICL. Its encoder extracts dense image features and projects them into key-value pairs stored in a dynamic memory. At query time, cross-attention over the keys assigns soft nearest-neighbor (NN) weights to aggregate values and predict novel views in-context. This modular design permits the use of any encoder.
In this work, we evaluate foundation model encoders for in-context object segmentation under unseen camera viewpoints. We extend the Hummingbird evaluation framework with the Multi-View ImageNet (MVImgNet) dataset [yu2023mvimgnetlargescaledatasetmultiview], grouping object views into angular bins. By constructing memory banks consisting of images from specific viewpoints and evaluating segmentation performance on held-out angles, we assess how well different encoders generalize across viewpoint shifts. Our main contribution combines three elements: (1) pixel-level segmentation, (2) extreme viewpoint changes, and (3) dynamic memory for ICL. Most existing work studies 3D understanding, but there is little evaluation of how consistent the features from frozen 2D foundation models remain when the camera moves to new viewpoints, and only a few in-context examples are available.
In addition to general-purpose ViT encoders [CLIP, DINO, DINOv2, DINOv3, SIGLIPv2, radiov2.5, tips_paper] we also include the Visual Geometry Grounded Transformer (VGGT) [VGGT], a model designed for 3D scene understanding. Since VGGT does not provide native patch embeddings, we rely on its aggregated tokens to construct a token sequence compatible with Hummingbird. This provides a basis for comparing contrastive pretraining approaches with a geometry-grounded multi-view model.
@@ 2 Related work
Recent developments in view generation have significantly improved the understanding of 3D scene structure in computer vision, setting new benchmarks along the way. A notable example is Neural Radiance Fields (NeRF) [mildenhall2020nerfrepresentingscenesneural], which achieves state-of-the-art novel view synthesis by optimizing a continuous volumetric function from sparse input images. While NeRF addresses generation, other studies investigate recognition in 3D understanding. ruan2023viewpointinvariantvisualrecognitionadversarial demonstrate that modern recognition models (ResNets [residual_learning], ViTs [dosovitskiy2020image], Swin [swin], and Masked Autoencoders (MAEs) [maskedAutoEncoders]) remain sensitive to viewpoint shifts in 3D scenes. With adversarial training, they improve invariance to 3D viewpoint changes beyond standard rotation-based augmentations. In parallel, shifman2024losttranslationmodernneural show that state-of-the-art encoders such as CLIP and DINOv2 change predictions under a single-pixel shift in up to 40% of cases, despite extensive augmentation during training. Although their analysis centers on 2D translation invariance, the findings highlight that robustness to geometric changes, whether in 2D or 3D, remains limited in current vision encoders.
@@
At a higher level, previous benchmarks in 3D computer vision were generally divided into two groups. The first group includes those working on traditional 3D recognition tasks, such as item detection (e.g., KITTI, Waymo [electronics14091882, Hwang2023Improvement]) or scene segmentation (e.g., SAI3D, ScanNet [Yin_2024_CVPR]), which typically test explicit 3D architectures or use classification/bounding box tasks. The second focuses on innovative view synthesis approaches, such as NeRF [mildenhall2020nerfrepresentingscenesneural], which prioritize dense creation rather than robust detection or segmentation under angular deviation. Beyond these 3D-focused benchmarks, general ICL benchmarks (such as the original Hummingbird framework [hummingbird] on PASCAL VOC [Everingham2010Jun]) assess feature transferability only in a 2D setting, and therefore fail to evaluate models under non-linear, wide-angle 3D viewpoint shifts.
@@
Beyond model architectures, multi-view datasets such as MVImgNet and PASCAL3D+ [Pascal3d] offer rich multi-view annotations that support the evaluation of cross-view generalization in object recognition. The rise of self-supervised Transformers has also opened new possibilities for downstream tasks such as segmentation. In particular, DINOv2 [DINO] produces robust embeddings that transfer effectively to prediction tasks, despite training without labels.
@@
Nevertheless, a gap remains across these benchmarks and models. Most prior work emphasizes detection, classification, or view-angle estimation, but not the measurement of 3D contextual understanding through multi-view segmentation. To address this, our benchmark primarily evaluates general-purpose foundation encoders, while also including a geometry-grounded model, VGGT [VGGT], to compare against approaches explicitly designed for 3D scene understanding. To our knowledge, no existing benchmark systematically evaluates a model’s ability to generalize segmentation across viewpoint shifts using dynamic memory. Our proposed evaluation method aims to address this gap in existing research.
@@ 3 Methodology
We evaluate the view generalization ability of frozen ViT models in semantic segmentation using a non-parametric, retrieval-based framework. Our approach builds on the Hummingbird framework [hummingbird], which applies ICL to vision tasks and supports the evaluation of spatial perception and semantic understanding. However, Hummingbird does not analyze performance variations across viewpoints changes. To address this, we introduce a viewpoint binning protocol and a multi-view dataset subset for cross-view robustness analysis. License information for the models, framework, and dataset can be found in Appendix A.
@@ 3.1 Models
We evaluate seven pretrained ViTs: CLIP [CLIP], SigLIP2 [SIGLIPv2], DINO [DINO], DINOv2 [DINOv2], DINOv3 [DINOv3], C-RADIOv2 [am-radio, radiov2.5], and TIPS [tips_paper]. Most models use a ViT-B/16 backbone, except DINOv2 and TIPS, which have a ViT-B/14 backbone. Our benchmark primarily focuses on general-purpose foundation encoders, but for comparison we also include the geometry-aware VGGT [VGGT]. This inclusion provides us a starting point to compare contrastive pretraining approaches and geometry-grounded representations.
@@ 3.2 Inference pipeline
We follow the Hummingbird [hummingbird] inference pipeline, where patch-level features from a frozen ViT are stored in a memory bank with one-hot semantic labels. During inference, query features are matched to the memory using FAISS with cosine similarity and $k=30$ nearest neighbors. The memory size is passed as an input parameter. Hummingbird’s cross-attention decoder aggregates the retrieved labels to produce segmentation predictions.
@@ 3.3 Dataset
Our study is based on MVImgNet, a large-scale dataset with over 6.5 million frames across 238 object categories [yu2023mvimgnetlargescaledatasetmultiview]. Each frame includes a segmentation mask, camera extrinsics, and a reconstructed 3D point cloud. This dataset is chosen for three reasons: (1) it contains various annotations across wide viewpoint ranges, enabling view generalization analysis; (2) it includes camera extrinsics that allow precise angular binning; and (3) it has scale and diversity, capturing a broad range of object appearances and spatial configurations.
@@ 3.3.1 Viewpoint binning
To study viewpoint robustness, we discretize relative camera angles into seven bins spanning 0–90, in steps of 15. To aid in this task, we use COLMAP [COLMAPsfm, COLMAPvote, COLMAPmvs]: a Structure-from-Motion (SfM) and Multi-View Stereo (MVS) pipeline that estimates camera poses and 3D scene geometry from a set of images. Using COLMAP extrinsics, we compute the relative rotation $R_{rel}$ between each frame and the first frame of the object instance. For each instance, we select one representative frame per bin by choosing the frame with the smallest angular error relative to the bin center. Images and masks are stored per bin for downstream use. The angular deviation $\theta$ is computed as
@@ 3.3.2 Subset construction
We curated a subset of MVImgNet, with the goal of selecting categories (classes) with sufficient angular coverage and manageable size for repeated memory-based evaluation. The two criteria used for selecting a category are: (1) the total zipped folder size must be 1–6 GB to ensure feasible data loading and storage; and (2) at least one object instance must span all seven angular bins with a maximum variance of 6. Categories that did not meet these criteria were excluded. For example, category 23 (laptop) was excluded, as it did not have an instance that reached an angle of 90. This yielded 15 segmentation classes, excluding the background. [^1] The 15 categories are ordered as follows: stove, sofa, microwave, bed, toy cat, toy cow, toy dragon, coat rack, guitar stand, ceiling lamp, toilet, sink, strings, broccoli and durian. More details can be found in Appendix B. Due to a misconfiguration in our codebase, eight classes had reduced per-bin image counts when evaluating the DINOv3 and VGGT models. All (complete and reduced) image counts per bin are found in Table 6 of Appendix B.
For each category instance, we parsed the COLMAP camera pose data to extract camera extrinsics and compute relative angles. For each valid instance, we selected the closest frame to each angular bin center and organized the resulting RGB images and masks in a structured directory grouped by category and angle. Additionally, we computed angular selection errors per bin (see Table 6). Examples of multi-view objects from the resulting dataset are shown in Figure 1.
Figure 1: Multi-view categories. Each MVImgNet category is shown across all viewpoint bins (0–90), above are example visualizations for 4 of the 15 selected classes.
@@ 3.4 Model input
To ensure a fair comparison, the largest native input size across the models is used: $504\times 504$ for models that use a patch size of 14 and $512\times 512$ for the ones using a patch size of 16. The evaluation of models with a bigger input size than their original resolution requires an additional step of embedding interpolation. In our setup, we upsample the absolute positional embeddings using bicubic interpolation to align with the new input resolution, as it provides smooth spatial transitions and is widely used for resizing in vision models. All evaluations are performed with a batch size of 4. The full configuration details are provided in Appendix C.
@@ 3.5 Metrics and multi-view evaluation
The segmentation quality is measured by mean Intersection over Union (mIoU) across all semantic classes. Each predicted segmentation mask consists of the background class and one of the 15 object classes. We report three metrics for each evaluation setting: (1) the per-class Intersection over Union (IoU), also known as the Jaccard index, (2) mIoU over all 16 classes, and (3) standard deviation. Results are presented per model, reference bin, and validation bin to analyze viewpoint robustness. Although all models perform well on the background class (see Figure 6), we include it in the mIoU computation for completeness.
@@ 4 Experimental results
To validate our setup, we first reproduced the Hummingbird evaluation on PASCAL VOC [Everingham2010Jun] using the official open-source implementation. [^2] The evaluation repository can be found at https://github.com/vpariza/open-hummingbird-eval. The results, shown in Table 9 of Appendix E.1, match the methodology and configuration of the original paper [hummingbird]. This confirms that our pipeline is consistent with the benchmark and provides a reliable basis for the experiments that follow.
With the setup validated, we proceed to evaluate viewpoint generalization and robustness across three experiments: (A) cross-viewpoint generalization, (B) breaking point analysis, and (C) memory size robustness.
4.1 Experiment A: cross-viewpoint generalization
We compare the models in terms of their ability to generalize across viewpoints. The goal is to assess how different pretraining strategies affect performance when only limited reference views are available in the memory. Specifically, we evaluate how well the models generalize to unseen viewpoints when trained on selected angular bins from the MVImgNet dataset. The reference and validation splits vary across four difficulty levels, which are summarized in Table 1. Each model is evaluated using the mIoU across all 16 classes (background and 15 object categories) and all validation images. This setup enables a controlled comparison of each model’s ability to interpolate between observed viewpoints and extrapolate to unseen ones.
Table 1: Difficulties and bin splits. Difficulty levels are defined by the choice of reference bins (stored in the memory bank) and validation bins (unseen views for testing generalization). An increasing difficulty corresponds to fewer reference bins and larger unseen angular gaps, which we expect makes interpolation and extrapolation progressively harder.
@@ 4.1.1 Results
As shown in Table 2, DINOv3 outperformed all models across all difficulties. DINO had the second-best performance for the Easy and Medium difficulty levels. Quantitative results illustrating these trends are shown in Figure 2, with per-class results in Appendix E.2. Across all classes, DINOv3, DINO, and DINOv2 had better segmentation performance under more difficult conditions than the other models. Overall, CLIP consistently followed close behind the DINO models, while C-RADIOv2, TIPS, and SigLIP2 performed noticeably worse, especially under limited reference view conditions. VGGT showed the lowest scores in our setting.
Table 2: mIoU scores across difficulty levels. For each model, we report the average mIoU and standard deviation over the four difficulty levels. The memory size used is 1,024k. DINOv3 achieves the best performance across all difficulties, followed by DINOv2 and DINO, of which the latter outperformed DINOv2 in the Easy and Medium setups (where multiple reference views are available). DINOv2 significantly outperformed DINO under the Extreme case with only a single reference bin. VGGT shows significantly worse results than other models.

### v2

@@ Abstract
Benchmarking 3D spatial understanding of foundation models is essential for real-world applications such as robotics and autonomous driving. Existing evaluations often rely on downstream fine-tuning with linear heads or task-specific decoders, making it difficult to isolate the intrinsic 3D reasoning ability of pre-trained encoders. In this work, we introduce a novel benchmark for in-context 3D scene understanding that requires no fine-tuning and directly probes the quality of dense visual features. Building on the Hummingbird framework, which evaluates in-context 2D scene understanding, we extend the setup to the 3D Multi-View ImageNet (MVImgNet) dataset. Given a set of images depicting objects at specific camera angles (keys), we benchmark the performance of segmenting novel views (queries) and report the scores in 4 categories of easy, medium, hard, and extreme based on the key-query view contrast. We benchmark 7 state-of-the-art foundation models and show that DINO-based encoders remain competitive across large viewpoint shifts. Our code is publicly available at https://github.com/ToyeshC/open-hummingbird-3d-eval.
@@ 1 Introduction
In recent years, we have seen the rise of foundation models [2, 28, 15, 32]. These large-scale pre-trained models support many tasks, including visual understanding, which enables direct comparison of their encoders. Despite their impressive capabilities, state-of-the-art Vision Transformers (ViTs) can suffer catastrophic failures when objects are viewed from unusual angles, as demonstrated by recent angle-sensitivity studies [20]. However, most vision benchmarks emphasize single-view tasks or dense synthesis [14, 17, 21, 22, 35], leaving segmentation robustness under camera rotations comparatively underexplored.
To address viewpoint variability, recent work has explored in-context learning (ICL) as a means to adapt models to unseen views without retraining. ICL is the ability of a model to perform new tasks by conditioning on a prompt [31]. An example is Hummingbird [1], a memory-augmented ViT for ICL. Its encoder extracts dense image features and projects them into key-value pairs stored in a dynamic memory. At query time, cross-attention over the keys assigns soft nearest-neighbor (NN) weights to aggregate values and predict novel views in-context. This modular design permits the use of any encoder.
In this work, we evaluate general-purpose ViT encoders [18, 3, 16, 27, 29, 8, 13] for in-context object segmentation under unseen camera viewpoints. We extend the Hummingbird evaluation framework with the Multi-View ImageNet (MVImgNet) dataset [34], grouping object views into angular bins. By constructing memory banks consisting of images from specific viewpoints and evaluating segmentation performance on held-out angles, we assess how well different encoders generalize across viewpoint shifts. Our main contribution combines three elements: (1) pixel-level segmentation, (2) extreme viewpoint changes, and (3) dynamic memory for ICL. Most existing work studies 3D understanding, but there is little evaluation of how consistent the features from frozen 2D foundation models remain when the camera moves to new viewpoints and only a few in-context examples are available.
@@ 2 Related work
Recent developments in view generation have significantly improved the understanding of 3D scene structure in computer vision, setting new benchmarks along the way. For example, Neural Radiance Fields (NeRFs) [14] achieve state-of-the-art novel view synthesis by optimizing a continuous volumetric function from sparse input images. While NeRFs addresses generation, other studies investigate recognition in 3D understanding. Ruan et al. [20] demonstrate that modern recognition models (ResNets [7], ViTs [4], Swin [12], and Masked Autoencoders (MAEs) [6]) remain sensitive to viewpoint shifts in 3D scenes. With adversarial training, they improve invariance to 3D viewpoint changes beyond standard rotation-based augmentations. In parallel, Shifman and Weiss [26] show that state-of-the-art encoders such as CLIP and DINOv2 change predictions under a single-pixel shift in up to 40% of cases, despite extensive augmentation during training. Although their analysis centers on 2D translation invariance, the findings highlight that robustness to geometric changes, whether in 2D or 3D, remains limited in current vision encoders.
@@
At a higher level, previous benchmarks in 3D computer vision were generally divided into two groups. The first group includes those working on traditional 3D recognition tasks, such as item detection (e.g., KITTI, Waymo [11, 9]) or scene segmentation (e.g., SAI3D, ScanNet [33]), which typically test explicit 3D architectures or employ classification/bounding box tasks. The second focuses on innovative view synthesis approaches, such as NeRF [14], which prioritize dense creation rather than robust detection or segmentation under angular deviation. Beyond these 3D-focused benchmarks, general ICL benchmarks such as the Hummingbird framework [1] or post-training approaches [22, 21, 17, 35] assess feature transferability only in a 2D setting, and therefore fail to evaluate models under non-linear, wide-angle 3D viewpoint shifts.
@@
Beyond model architectures, multi-view datasets such as MVImgNet and PASCAL3D+ [30] offer rich multi-view annotations that support the evaluation of cross-view generalization in object recognition. The rise of self-supervised Transformers has also opened new possibilities for downstream tasks such as segmentation. In particular, DINOv2 [3] produces robust embeddings that transfer effectively to prediction tasks, despite training without labels. More recently, Siméoni et al. [27] introduced DINOv3, which significantly scales the DINOv2 framework. A critical technical contribution of DINOv3 is the introduction of Gram Anchoring, a mechanism designed to prevent the degradation of dense feature maps during long-duration self-supervised training. By regularizing patch-level similarity structures against a teacher model, this model maintains high-quality spatial features that are crucial for dense prediction tasks.
@@
Nevertheless, a gap remains across these benchmarks and models. Most prior work emphasizes detection, classification, or view-angle estimation, but not the measurement of 3D contextual understanding through multi-view segmentation. To address this, our benchmark primarily evaluates general-purpose foundation encoders. To our knowledge, no existing benchmark systematically evaluates a model’s ability to generalize segmentation across viewpoint shifts using dynamic memory. Our proposed evaluation method aims to address this gap in existing research.
@@ 3 Methodology
We evaluate the view generalization ability of frozen ViT models in semantic segmentation using a non-parametric, retrieval-based framework. Our approach builds on the Hummingbird framework [1], which applies ICL to vision tasks and supports the evaluation of spatial perception and semantic understanding. However, Hummingbird does not analyze performance variations across viewpoints changes. To address this, we introduce a viewpoint binning protocol and a multi-view dataset subset for cross-view robustness analysis. License information for the models, framework, and dataset can be found in Appendix A.
@@ 3.1 Models
We evaluate 7 pre-trained ViTs: CLIP [18], SigLIP2 [29], DINO [3], DINOv2 [16], DINOv3 [27], C-RADIOv2 [19, 8], and TIPS [13]. Most models use a ViT-B/16 backbone, except DINOv2 and TIPS, which have a ViT-B/14 backbone.
@@ 3.2 Inference pipeline
We follow the Hummingbird [1] inference pipeline, where patch-level features from a frozen ViT are stored in a memory bank with one-hot semantic labels. During inference, query features are matched to the memory using FAISS [^1] Facebook AI Similarity Search (FAISS) is an open-source library designed for efficient similarity search and clustering of dense vectors, optimized for high-dimensional feature retrieval [10]. with cosine similarity and $k=30$ nearest neighbors. The memory size is passed as an input parameter. Hummingbird’s cross-attention decoder aggregates the retrieved labels to produce segmentation predictions.
@@ 3.3 Dataset
Our study is based on MVImgNet, a large-scale dataset with over 6.5 million frames across 238 object categories [34]. Each frame includes a segmentation mask, camera extrinsics, and a reconstructed 3D point cloud. This dataset is chosen for three reasons: (1) it contains various annotations across wide viewpoint ranges, enabling view generalization analysis; (2) it includes camera extrinsics that allow precise angular binning; and (3) it has scale and diversity, capturing a broad range of object appearances and spatial configurations.
@@ 3.3.1 Viewpoint binning
To study viewpoint robustness, we discretize relative camera angles into 7 bins spanning 0–90, in steps of 15. To aid in this task, we use COLMAP [23, 24, 25]: a Structure-from-Motion (SfM) and Multi-View Stereo (MVS) pipeline that estimates camera poses and 3D scene geometry from a set of images. Using COLMAP extrinsics, we compute the relative rotation $R_{rel}$ between each frame and the first frame of the object instance. For each instance, we select 1 representative frame per bin by choosing the frame with the smallest angular error relative to the bin center. Images and masks are stored per bin for downstream use. The angular deviation $\theta$ is computed as
@@ 3.3.2 Subset construction
We curated a subset of MVImgNet, with the goal of selecting categories (classes) with sufficient angular coverage and manageable size for repeated memory-based evaluation. The two criteria used for selecting a category are: (1) the total zipped folder size must be 1–6 GB to ensure feasible data loading and storage; and (2) at least one object instance must span all 7 angular bins with a maximum variance of 6. Categories that did not meet these criteria were excluded. For example, category 23 (laptop) was excluded, as it did not have an instance that reached an angle of 90. This yielded 15 segmentation classes, excluding the background. [^2] The 15 categories are ordered as follows: stove, sofa, microwave, bed, toy cat, toy cow, toy dragon, coat rack, guitar stand, ceiling lamp, toilet, sink, strings, broccoli, and durian. More details can be found in Appendix B.
For each category instance, we parsed the COLMAP camera pose data to extract camera extrinsics and compute relative angles. For each valid instance, we selected the closest frame to each angular bin center and organized the resulting RGB images and masks in a structured directory grouped by category and angle. Additionally, we computed angular selection errors per bin (see Table 6). Examples of different views of several class instances from the resulting dataset are shown in Figure 1.
Figure 1: Multi-view instances. Each MVImgNet category contains instances spanning all discretized viewpoint bins between 0 and 90. The figure shows example visualizations of 4 instances from the 15 selected classes.
@@ 3.4 Model input
To ensure a fair comparison, we use the largest native input size across all models. This is $504\times 504$ for models that use a patch size of 14 and $512\times 512$ for the ones using a patch size of 16. When evaluating models at resolutions higher than those used during pre-training, we resize the pre-trained absolute positional embeddings to match the new patch grid using bicubic interpolation. All evaluations are performed with a batch size of 4. The full configuration details are provided in Appendix C.
@@ 3.5 Metrics and multi-view evaluation
The segmentation quality is measured by mean Intersection over Union (mIoU) across all semantic classes. Each predicted segmentation mask consists of the background class and one of the 15 object classes. We report 3 metrics for each evaluation setting: the per-class Intersection over Union (IoU), also known as the Jaccard index, mIoU over all 16 classes, and standard deviation. Results are presented per model, reference bin, and validation bin to analyze viewpoint robustness. Although all models perform well on the background class (see Figure 6), we include it in the mIoU computation for completeness.
@@ 4 Experimental results
To validate our setup, we first reproduced the Hummingbird evaluation on PASCAL VOC [5] using the official open-source implementation. [^3] The evaluation repository can be found at https://github.com/vpariza/open-hummingbird-eval. The results, shown in Table 8 of Appendix E, match the methodology and configuration of the original paper [1]. This confirms that our pipeline is consistent with the benchmark and provides a reliable basis for the experiments that follow.
With the setup validated, we proceed to evaluate viewpoint generalization and robustness across 3 experiments: (A) cross-viewpoint generalization, (B) breaking point analysis, and (C) memory size robustness.
Table 1: Difficulties and bin splits. Difficulty levels are defined by the choice of reference bins (stored in the memory bank) and validation bins (unseen views for testing generalization). An increasing difficulty corresponds to fewer reference bins and larger unseen angular gaps, which makes interpolation and extrapolation progressively harder.
@@
4.1 Experiment A: cross-viewpoint generalization
We compare the models in terms of their ability to generalize across viewpoints. The goal is to assess how different pre-training strategies affect performance when only limited reference views are available in the memory. Specifically, we evaluate how well the models generalize to unseen viewpoints when trained on selected angular bins from the MVImgNet dataset. The reference and validation splits vary across 4 difficulty levels, which are summarized in Table 1. Each model is evaluated using the mIoU across all 16 classes (background and 15 object categories) and all validation images. This setup enables a controlled comparison of each model’s ability to interpolate between observed viewpoints and extrapolate to unseen ones.
@@ 4.1.1 Results
As shown in Table 2, DINOv3 outperformed all models across all difficulties. DINO had the second-best performance for the Easy and Medium difficulty levels. Quantitative results illustrating these trends are shown in Figure 2, with per-class results in Appendix F.1. Across all classes, DINOv3 and DINOv2 maintain stronger segmentation performance as difficulty increases, with DINO remaining competitive under most settings. CLIP performs close to the DINO models and surpasses DINO under the Extreme setting, while C-RADIOv2, TIPS, and SigLIP2 perform noticeably worse, particularly when reference views are limited.
Table 2: mIoU scores across difficulty levels. For each model, we report the average mIoU and standard deviation over the 4 difficulty levels. The memory size used is 1,024k. DINOv3 achieves the best performance across all difficulties, followed by DINOv2 and DINO, of which the latter outperformed DINOv2 in the Easy and Medium setups (where multiple reference views are available). DINOv2 significantly outperformed DINO under the Extreme case with only a single reference bin.